Mary is 3 years older than tim and twice as old as anna. Tim is 2 years older than anna. What are the ages of mary, tim, and anna?
Answer by ankor@dixie-net.com(22740)

Mary is 3 years older than tim and twice as old as anna. Tim is 2 years older than anna. What are the ages of mary, tim, and anna
:
Write an equation for each statement:
Mary is 3 years older than tim
m = t + 3
:
and twice as old as anna.
m = 2a
:
Tim is 2 years older than anna.
t = a + 2
:
Using the 1st equation, replace t with (a+2), replace m with 2a
2a = (a+2) + 3
2a = a + 5
2a - a = 5
a = 5 yrs is Anna's age
then
t = 5 + 2
t = 7 yrs is Tim's age
and
m = 2(5)
m = 10 yrs is Mary's age
:
These are the ages of mary, tim, and anna